solve:
5(x-2)^2=3

Answer by malakumar_kos@yahoo.com(315)


solve:
5(x-2)^2=3 consider 5(x-2)^2 = 3
5(x^2-4x+4)-3 = 0
5x^2-20x+20-3 = 0
5x^2-20x+17 = 0
use the standard formula x = -b+-sq rtb^2-4ac/2a
where a = 5 b = -20 c = 17
x = -(-20)+-sq rt(-20)^2-4(5)(17)/2.5
= 40+-sq rt(400-340)/10
= 40+-sq rt60/10
= 40+-2sq rt15/10
= 2(20+-sq rt15)/10
= (20+-sq rt15)/5 is the required solution